# Towards Exploring Continual Fine-Tuning for Enhancing Language Ability in Large Language Models

### v1

@@ 3.2 Task and Language Ability Results
Model Phase 1 Phase 2 MLQA ()($\uparrow$)( ↑ ) XLSUM ()($\uparrow$)( ↑ ) XQuAD ()($\uparrow$)( ↑ ) Average Dataset Dataset Phase 1 Phase 2 Phase 1 Phase 2 Phase 1 Phase 2 Phase 1 Phase 2 Mistral-7B Alpaca MultiAlpaca 0.229 0.288 0.012 0.060 0.290 0.602 0.177 0.317 Instruct 0.246 0.307 0.012 0.033 0.351 0.436 0.203 0.259 LLaMA-3-8B Alpaca 0.438 0.597 0.033 0.034 0.586 0.737 0.352 0.456 Instruct 0.609 0.321 0.048 0.027 0.712 0.417 0.456 0.255
@@
Table 2: Language Ability results for two-phase Continual Fine-tuning (CFT). With green, we denote an improvement in language ability post Phase 2 fine-tuning. Likewise, we denote a decline in language ability with red. For MLQA and XQUAD we use F1 abstractive score, while for XLSUM we use ROUGE Score.
@@
We compare the task and language ability of Mistral-7B and LLaMA-3-8B continually fine-tuned models on different phase-wise datasets [^4] When it is clear from the context, we use “Instruct” to denote the dataset used in Phase 1 to instruction fine-tune Mistral-7B-Instruct or LLaMA-3-8B-Instruct.. Table 1 presents the results for task ability, while Table 2 presents the results for language ability. Table 2 reports the average score across languages. We provide language-specific scores in the Appendix (§B).
@@
Results Discussion. From Table 1, we see that for phase-wise datasets like Instruct and MultiAlpaca, the performance of the Phase 2 models trained on them declines for English. This decline occurs when they are continually fine-tuned on multilingual data in Phase 2. However, we see a jump in Mistral-7B’s language ability from the results for the multilingual generative tasks (Table 2). These models fine-tuned on multilingual datasets show catastrophic forgetting in English. However, for phase-wise datasets like Alpaca followed by MultiAlpaca, we see that models trained on them do not show a decline in task ability (Table 1). We also see a gain in these models’ language ability (Table 2) [^5] LLaMA-3-8B Instruct MultiAlpaca shows deterioration in LA. We explain this behavior in §4.3..
@@
Additional Ablations. In the Appendix (§B), we also present results for OpenOrca-mOpenOrca phase-wise datasets. For Mistral-7B, we observe that the average task ability of the Phase 2 model (over Phase 1’s Mistral-7B-OpenOrca) marginally declines: 0.487 from 0.504. Whereas, for Mistral-7B-Instruct, the average decline in task ability is significant: 0.376 from 0.529. Likewise, for LLaMA-3-8B, the average task ability for LLaMA-3-8B OpenOrca mOpenOrca sees an increase of 0.415 from 0.404. In contrast, with Instruct-mOpenOrca as the phase-wise datasets, the task ability significantly drops, from 0.437 to 0.173.
@@
Observation. With Table 1, we see that our two-phase CFT setup for language adaption shows an interesting trend: for certain pairs of phase-wise datasets (e.g., Alpaca & MultiAlpaca), the LLM after Phase 2 sees an improvement in the task ability (computed on English evaluation tasks). We notice that phase-wise datasets like Alpaca and MultiAlpaca have the same seed prompts. Alternately, the two datasets encode the same tasks in different languages. We hypothesize an LLM fine-tuned on either of these datasets learns the same task ability, and therefore, the second phase of CFT leads to lesser interference in the representation space. That is, an LLM continually fine-tuned on Alpaca & MultiAlpaca preserves its task ability across phases. We next define two metrics that aim to quantify the task-specific similarity of two datasets.
3.3 Phase-wise Datasets: Similarity of Representations
@@
To quantify whether two datasets encode the same tasks, we define DES that computes a similarity score using the dot product of the average representations (embedding) generated by a language-agnostic model.
@@
$f_{\textsf{DES}}(D_{1},D_{2};\Theta)=\langle\mathbf{E}_{\Theta}(D_{1}),\mathbf{E}_{\Theta}(D_{2})\rangle$ (1)
@@
Higher the DES score, more similar the embedding, i.e., greater similarity between $D_{1}$ and $D_{2}$. For $\Theta$, we use the language-agnostic sentence-tokenizer LaBSE (feng2020languageagnostic). We compute DES by encoding 500 random samples from Alpaca, MultiAlpaca, OpenOrca, and mOpenOrca, and measure $f_{\textsf{DES}}$ for each pair.
Fixing Alpaca as the Phase 1 dataset $D_{1}$, when the Phase 2 dataset $D_{2}$ is MultiAlpaca, the DES score is 0.924 and 0.792 for mOpenOrca. When $D_{1}$ is OpenOrca, the DES score for mOpenOrca as $D_{2}$ is 0.953 and 0.774 when $D_{2}$ is MultiAlpaca. For dataset pairs with similar tasks, we see a high DES score and relatively low scores for datasets with different tasks. That is, DES captures the (pair-wise) variation in task abilities of these datasets.
Model Parameter Difference (MPD). Another method of quantifying the similarity of tasks for two datasets $D_{1}$ and $D_{2}$ is to compute the difference between the parameters of models $\Theta_{1}$ (fine-tuned on $D_{1}$) and $\Theta_{2}$ (fine-tuned on $D_{2}$). Geometrically, the difference of the parameters captures the representation shift of $\Theta_{2}$ in the space defined by $\Theta_{1}$. If $D_{1}$ $D_{2}$ encode the same tasks, the combined shift by $\Theta_{2}$ should be relatively lower, compared to the shift if $D_{1}$ $D_{2}$ encode different tasks. Formally,
@@
$f_{\textsf{MPD}}(\Theta_{1},\Theta_{2};\Theta_{B})=\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{w}(\Theta_{1,i})-\mathbf{w}(\Theta_{2,i})\|_{2}$ (2)
[…]
The smaller the MPD score, the closer the fine-tuned models are in the parameter space. Fixing Mistral-7B as the base model $\Theta_{B}$, and $D_{1}$ as MultiAlpaca, we vary $D_{2}$ as one of AlpacaOpenOrca, and mOpenOrca, and observe the corresponding MPD scores. We normalize the MPD scores with the maximum observed score across all three models for a fair comparison. With $D_{2}$ as Alpaca, the MPD score is $0.294$. For $D_{2}$ as Instruct, MPD is $1.0$ and 0.55 when $D_{2}$ is OpenOrca. These scores show a similar trend to DES: for Alpacaand MultiAlpaca, the scores are lower, highlighting the similarities in the datasets in the parameter space. We see relatively higher scores for the other pair of models, implying a difference in the dataset pairs.
3.4 Visualizing Decline in Task Ability
@@
To explain the effect of similar phase-wise data sets on the LLM task ability, we look at the model representations when parsing English (as the task ability is computed over English). We feed MTBench (zheng2024judging) to the models, an English prompt dataset for testing, and visualize the similarity between the mean hidden activations, for each model layer. For the analysis, given an LLM $\Theta$ with $l$ layers, let $X_{\Theta}\in\mathbb{R}^{l\times d}$ be the mean hidden activations, across $n$ samples from MTBench.
Figure 2: We see a greater change in the variation of the representations for non-similar datasets (e.g., Instruct & MultiAlpaca) compared to similar datasets (e.g., Alpaca & MultiAlpaca). Interestingly, for LLaMA-3-8B the change is large across layers and a magnitude higher than Mistral-7B. For Mistral-7B, we see the later layers showing the most change.
t-SNE Visualization. Figure 1 depicts t-SNEs (vanDerMaaten2008) for $X_{\textsc{Mistral-7B}}$ and $X_{\textsc{LLaMA-3-8B}}$ LLMs, continually fine-tuned on the phase-wise datasets Alpaca & MultiAlpaca and Instruct & MultiAlpaca. We observe that for similar phase-wise datasets, the model before and after Phase 2 produces similar hidden activations. Contrarily, for non-similar phase-wise datasets, the hidden activations form distinct clusters, implying separation between the phase-wise activations. That is, the model representations for non-similar phase-wise datasets are well-separated. The model representations during Phase 2 do not align with Phase 1 representations; thus, resulting in greater model weight interference leading to a decline in task ability.
Visualizing Variance in Model Representations. Figure 1 provides some intuition for the correlation between phase-wise datasets and decline in task ability. To further understand the layer-wise behavior of the hidden activations, similar to chang-etal-2022-geometry, we compute covariance matrices $\Sigma_{\Theta}$ for each $X_{\Theta}$. Intuitively, $\Sigma_{\Theta}$ captures the variance in different directions for representations of hidden activations for $\Theta$.
We first compute the mean centered activation matrix $\bar{X}_{\Theta}=X_{\Theta}-\mu_{\Theta}$, according to $\mu_{\Theta}\in\mathbb{R}^{d}$. Next, we derive $\Sigma_{\Theta}=\frac{1}{l-1}\cdot\bar{X}_{\Theta}^{T}\bar{X}_{\Theta}\in\mathbb{R}^{d\times d}$. To compare the layer-wise variance in representations, we compute the L2-Norm of the difference of the matrices $\Sigma_{\textsc{Mistral-7B}}$ (Figure 2 (left)) or $\Sigma_{\textsc{LLaMA-3-8B}}$ (Figure 2 (right)) when continually fine-tuned on Alpaca & MultiAlpaca (blue lines) or Instruct & MultiAlpaca (red lines).
@@
CFT Setup Task Ability (TA) Language Ability (LA) Phase 2 Mitigating IFEval Alpaca Eval MMLU HellaSwag Avg MLQA XLSum XQUAD Avg Dataset Strategy ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) Mistral-7B 0.55 0.35 0.575 0.641 0.529 0.246 0.012 0.351 0.203 MultiAlpaca 0.462 0.15 0.533 0.416 0.390 0.307 0.033 0.436 0.259 LF_H1 0.456 0.03 0.497 0.598 0.395 0.176 0.016 0.215 0.136 LF_H2 0.364 0.12 0.364 0.504 0.338 0.213 0.014 0.442 0.223 GR_5 0.540 0.17 0.540 0.611 0.465 0.311 0.008 0.428 0.249 GR_10 0.567 0.12 0.567 0.594 0.462 0.213 0.007 0.427 0.215 LoRA 0.383 0.09 0.579 0.625 0.42 0.289 0.043 0.518 0.283 ER_10 0.593 0.08 0.580 0.635 0.599 0.249 0.008 0.398 0.218 LLaMA-3-8B 0.735 0.14 0.340 0.533 0.436 0.609 0.048 0.712 0.456 MultiAlpaca 0.182 0.10 0.239 0.278 0.217 0.321 0.030 0.417 0.256 LF_H1 0.303 0.0 0.231 0.275 0.202 0.368 0.037 0.505 0.303 LF_H2 0.380 0.06 0.485 0.525 0.373 0.400 0.038 0.505 0.314 GR_5 0.269 0.01 0.516 0.316 0.279 0.437 0.019 0.593 0.349 GR_10 0.264 0.12 0.229 0.250 0.228 0.254 0.009 0.314 0.192 LoRA 0.196 0.0 0.280 0.235 0.179 0.007 0.008 0.005 0.007 ER_10 0.420 0.02 0.603 0.561 0.420 0.434 0.025 0.53 0.330
@@
Table 3: Task and Language ability results for our mitigating strategies, Generative Replay (GR_5 & GR_10) and Layer Freezing (LF_H1 & LF_H2). We also use LoRA (hu2022lora) and ER_10 as two baseline strategies. Here, we perform Phase 2 fine-tuning with rank 64 and quantisation bfloat16 for LoRA. For ER_10, we use the English dataset used in GR_5 with original responses. The Phase 1 dataset is Instruct for each row. The first two rows for both Mistral-7B and LLaMA-3-8B provide numbers for Instruct and Instruct-MultiAlpaca (from Table 1 & Table 2).
4 Mitigating Strategies for Multilingual CFT

### v3

@@ 4.2 Results
Two-phase Continual Fine-tuning Model Phase 1 (P1) Phase 2 (P2) MLQA $(\uparrow)$ XLSUM $(\uparrow)$ XQuAD $(\uparrow)$ GMMLU $(\uparrow)$ Average Dataset Dataset P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 Mistral-7B Alpaca 0.229 0.288 0.012 0.060 0.290 0.602 0.42 0.40 0.238 0.338 Instruct Multi 0.246 0.307 0.012 0.033 0.351 0.436 0.44 0.43 0.262 0.302 LLaMA-3-8B Alpaca Alpaca 0.438 0.597 0.033 0.034 0.586 0.737 0.53 0.34 0.397 0.427 Instruct 0.609 0.321 0.048 0.027 0.712 0.417 0.25 0.44 0.405 0.301 Dataset Mixture Model Dataset Mixture MLQA $(\uparrow)$ XLSUM $(\uparrow)$ XQuAD $(\uparrow)$ GMMLU $(\uparrow)$ Average Mistral-7B Alpaca MultiAlpaca 0.406 0.079 0.217 0.41 0.278 LLaMA-3-8B Alpaca MultiAlpaca 0.480 0.040 0.139 0.50 0.289
@@
Table 2: Multilingual Ability results for two-phase Continual Fine-tuning (CFT). With green, we denote an improvement in Multilingual Ability post Phase 2 fine-tuning. Likewise, we denote a decline in Multilingual Ability with red. For MLQA and XQUAD we use F1 abstractive score, while for XLSUM we use ROUGE Lin (2004) score. We also provide numbers for dataset mixture – when the models are fine-tuned simultaneously on the Phase 1 and Phase 2 datasets.
@@
We compare the English and Multilingual ability of Mistral-7B and LLaMA-3-8B continually fine-tuned models on different phase-wise datasets [^5] When it is clear from the context, we use “Instruct” to denote the dataset used in Phase 1 to instruction fine-tune Mistral-7B-Instruct or LLaMA-3-8B-Instruct.. Table 1 presents the results for English Ability (EA), while Table 2 presents the results for Multilingual Ability (MA). Table 2 reports the average score across languages. We provide language-specific scores and results when the phases are reversed (e.g., MultiAlpaca-Alpaca) in §B and §C.
Comparison with Mixture.
From Tables 1 & 2: for Mixture, the mean of EA and MA scores for Mistral-7B fine-tuned on Alpaca-MultiAlpaca is 0.325, and 0.312 for LLaMA-3-8B. The corresponding two-phase mean score is 0.355 for Mistral-7B and 0.346 for LLaMA-3-8B. Our two-phase CFT setup is more effective than Mixture, for approximately the same number of training steps.
Discussion.
@@
From Table 1, for phase-wise datasets like Instruct and MultiAlpaca, the performance of the Phase 2 models trained on them declines for English. This decline occurs when they are continually fine-tuned on multilingual data in Phase 2. However, we see a jump in Mistral-7B’s multilingual ability for the multilingual generative tasks (Table 2). That is, Phase 2 models fine-tuned on multilingual datasets show forgetting in English. However, for phase-wise datasets like Alpaca followed by MultiAlpaca, we see that Phase 2 models do not show a decline in English ability (Table 1). We also see a gain in these models’ multilingual ability (Table 2).
Ablations.
@@
In Tables B2 & B3 (§B), we present results for OpenOrca-mOpenOrca phase-wise datasets. First, the "dataset mixture" again performs worse on average than CFT: 0.186 vs. 0.372 for Mistral-7B and 0.217 vs. 0.366 for LLaMA-3-8B. Second, for Mistral-7B, the average English ability of the Phase 2 model (over Phase 1’s Mistral-7B-OpenOrca) marginally improves: 0.413 from 0.407. Whereas, for Mistral-7B-Instruct, the average decline in English ability is significant: 0.30 from 0.450. Likewise, for LLaMA-3-8B, the average English ability for LLaMA-3-8B OpenOrca mOpenOrca sees an increase to 0.356 from 0.337. In contrast, for Instruct-mOpenOrca, the English ability significantly drops, from 0.372 to 0.138.
Observation.
@@
With Table 1, we see that our two-phase CFT setup for multilingual adaptation shows an interesting trend: for certain pairs of phase-wise datasets (e.g., Alpaca & MultiAlpaca), the LLM after Phase 2 sees an improvement in the English ability (computed on English evaluation tasks). We notice that phase-wise datasets like Alpaca and MultiAlpaca have the same seed prompts. Alternately, the two datasets encode the same instructions in different languages. We hypothesize an LLM fine-tuned on either of these datasets learns the same instructions, and therefore, the second phase of CFT leads to lesser interference in the representation space. That is, an LLM continually fine-tuned on Alpaca & MultiAlpaca preserves its English ability across phases. We next define two metrics that aim to quantify the instruction-specific similarity of two datasets.
4.3 Similarity of Phase-wise Datasets
Phase 1 Dataset Phase 2 Dataset DES ($\uparrow$) Alpaca MultiAlpaca 0.924 mOpenOrca 0.792 OpenOrca mOpenOrca 0.953 MultiAlpaca 0.774 Mistral-7B Instruct MultiAlpaca 0.746 : Prepared using model responses on MTBench Zheng et al. (2024)
Table 3: Quantifying Phase-wise Dataset Similarity using DES: higher the score, greater the dataset similarity.
@@ 4.3 Similarity of Phase-wise Datasets
To quantify whether two datasets are similar [^6] The CL-ML literature often defines task similarity via permutation tasks, emphasizing input-output transformations Goldfarb et al. (2024). Whereas, we consider semantic and structural similarity in natural language instructions., we define DES that computes a similarity score using the dot product of the average representations (embeddings) generated by a language-agnostic model.
@@
$f_{\textsf{DES}}(D_{1},D_{2};\Theta)=\langle\mathbf{E}_{\Theta}(D_{1}),\mathbf{E}_{\Theta}(D_{2})\rangle$
@@
Higher the DES score, more similar the embedding, i.e., greater similarity between $D_{1}$ and $D_{2}$. For $\Theta$, we use the language-agnostic sentence-tokenizer LaBSE Feng et al. (2020). We compute DES by encoding 500 random samples from Alpaca, MultiAlpaca, OpenOrca, and mOpenOrca, and measure $f_{\textsf{DES}}$ for each pair. Table 3 presents the numbers. For dataset pairs with similar datasets, we see a high DES score and relatively low scores for dissimilar datasets. DES captures the (pair-wise) variation in instruction similarity of these datasets.
Dataset $D_{2}$ Model Parameter Difference ($\downarrow$) Alpaca 0.29 Instruct 1.00 OpenOrca 0.55
Table 4: Quantifying Phase-wise Similarity using MPD: lower the score, greater the dataset similarity. Here, we fix MultiAlpaca as $D_{1}$ and $\theta_{B}$ as Mistral-7B.
Model Parameter Difference (MPD).
Another method of quantifying the similarity of instructions for two datasets $D_{1}$ and $D_{2}$ is to compute the difference between the parameters of models $\Theta_{1}$ (fine-tuned on $D_{1}$) and $\Theta_{2}$ (fine-tuned on $D_{2}$). Geometrically, the difference of the parameters captures the representation shift of $\Theta_{2}$ in the space defined by $\Theta_{1}$. If $D_{1}$ & $D_{2}$ encode the same datasets, the combined shift by $\Theta_{2}$ should be relatively lower, compared to the shift if $D_{1}$ & $D_{2}$ encode different intstructions. Formally,
@@
$f_{\textsf{MPD}}(\Theta_{1},\Theta_{2};\Theta_{B})=\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{w}(\Theta_{1,i})-\mathbf{w}(\Theta_{2,i})\|_{2}$
Here, $\mathbf{w}(\Theta_{j,i}),~\forall j\in\{1,2\}$ is $\Theta_{j}$’s $i\textsuperscript{th}$ parameter.
The smaller the MPD score, the closer the fine-tuned models are in the parameter space. Fixing Mistral-7B as the base model $\Theta_{B}$, and $D_{1}$ as MultiAlpaca, we vary $D_{2}$ as one of AlpacaOpenOrca, and mOpenOrca, and observe the corresponding MPD scores. We normalize the MPD scores with the maximum observed score across all three models for a fair comparison (see Table 4). MPD shows a similar trend to DES: for Alpaca and MultiAlpaca, the scores are lower, highlighting the similarities in the datasets in the parameter space. We see relatively higher scores for the other pair of models, implying a difference in the dataset pairs.
4.4 Visualizing Decline in English Ability
@@
To explain the effect of similar phase-wise data sets on an LLM’s EA, we look at model representations when parsing English. We feed MTBench Zheng et al. (2024) to the models, a widely-used English benchmark for generalized instruction-following evaluation, and visualize the similarity between the mean hidden activations for each model layer. For the analysis, given an LLM $\Theta$ with $l$ layers, let $X_{\Theta}\in\mathbb{R}^{l\times d}$ be the mean hidden activations, across $n$ samples from MTBench.
t-SNE Visualization.
Figure 1 depicts t-SNEs van der Maaten and Hinton (2008) for $X_{\textsc{Mistral-7B}}$ and $X_{\textsc{LLaMA-3-8B}}$ when these are continually fine-tuned on (i) Alpaca & MultiAlpaca and (ii) Instruct & MultiAlpaca. We observe that for similar phase-wise datasets, the model before and after Phase 2 produces similar hidden activations. Contrarily, for non-similar phase-wise datasets, the hidden activations form distinct clusters, implying separation between the phase-wise activations. That is, the model representations for non-similar phase-wise datasets are well-separated. The separation between model representations results in increased weight interference during Phase 2 – leading to a decline in EA.
Visualizing Variance in Model Representations.
Figure 1 provides an intuition for the correlation between phase-wise datasets and the decline in English ability. To further understand the layer-wise behavior of the hidden activations, similar to Chang et al. (2022), we compute covariance matrices $\Sigma_{\Theta}$ for each $X_{\Theta}$. Intuitively, $\Sigma_{\Theta}$ captures the variance in different directions for representations of hidden activations for $\Theta$.
Figure 2: Visualizing variance in model representations. We plot $\|\Sigma_{\text{Phase 2}}-\Sigma_{\text{Phase 1}}\|_{2}$ across layers to measure representation shift. Dissimilar datasets (e.g., Instruct & MultiAlpaca) induce substantially larger variation than similar ones (e.g., Alpaca& MultiAlpaca). For LLaMA-3-8B, this variation is consistently high across layers, whereas for Mistral-7B, it is more localized and increases primarily in higher layers.
We first compute the mean centered activation matrix $\bar{X}_{\Theta}=X_{\Theta}-\mu_{\Theta}$, where $\mu_{\Theta}=1/l\sum_{i=1}^{l}X_{\Theta}^{(i)}$. Next, we derive $\Sigma_{\Theta}=\frac{1}{l-1}\cdot\bar{X}_{\Theta}^{T}\bar{X}_{\Theta}\in\mathbb{R}^{d\times d}$. To compare the layer-wise variance in representations, we compute the L2-Norm of the difference of the matrices $\Sigma_{\textsc{Mistral-7B}}$ (Figure 2 (left)) or $\Sigma_{\textsc{LLaMA-3-8B}}$ (Figure 2 (right)) when continually fine-tuned on Alpaca & MultiAlpaca (blue lines) or Instruct & MultiAlpaca (red lines).
@@
CFT Setup English Ability (EA) Multilingual Ability (MA) Combined Mitigating IFEval Alpaca Eval MMLU HellaSwag XLSUM_en Avg MLQA XLSUM XQUAD GMMLU Avg Avg Strategy $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ Mistral-7B 0.462 0.15 0.533 0.416 0.10 0.332 0.307 0.033 0.436 0.430 0.302 0.317 LF_H1 0.456 0.03 0.497 0.598 0.10 0.336 0.176 0.016 0.215 0.428 0.209 0.273 LF_H2 0.364 0.12 0.364 0.504 0.12 0.294 0.213 0.014 0.442 0.384 0.263 0.279 Spectrum 0.435 0.24 0.488 0.524 0.13 0.363 0.317 0.083 0.176 0.370 0.237 0.30 GR_5 0.540 0.17 0.540 0.611 0.11 0.394 0.311 0.008 0.428 0.445 0.298 0.346 GR_10 0.567 0.12 0.567 0.594 0.12 0.394 0.213 0.007 0.427 0.450 0.274 0.334 ER_10 0.593 0.08 0.580 0.635 0.13 0.404 0.249 0.008 0.398 0.448 0.276 0.340 LoRA 0.383 0.09 0.579 0.625 0.03 0.341 0.289 0.043 0.518 0.435 0.321 0.331 LLaMA-3-8B 0.182 0.10 0.239 0.278 0.09 0.178 0.321 0.030 0.417 0.440 0.302 0.240 LF_H1 0.303 0.0 0.231 0.275 0.01 0.164 0.368 0.037 0.505 0.237 0.287 0.225 LF_H2 0.380 0.06 0.485 0.525 0.08 0.306 0.400 0.038 0.505 0.338 0.320 0.313 Spectrum 0.409 0.09 0.612 0.524 0.01 0.329 0.429 0.056 0.086 0.472 0.261 0.295 GR_5 0.269 0.01 0.516 0.316 0.07 0.236 0.437 0.019 0.593 0.342 0.348 0.292 GR_10 0.264 0.12 0.229 0.250 0.0 0.173 0.254 0.009 0.314 0.238 0.204 0.189 ER_10 0.420 0.02 0.603 0.561 0.12 0.345 0.434 0.025 0.53 0.448 0.359 0.352 LoRA 0.196 0.0 0.280 0.235 0.0 0.142 0.007 0.008 0.005 0.278 0.075 0.109
@@
Table 5: English Ability (EA) and Multilingual Ability (MA) results for our mitigating strategies. These comprise Generative Replay (GR_5 & GR_10), English Replay (ER_10) and Layer Freezing (LF_H1, LF_H2 & Spectrum). We use LoRA Hu et al. (2022) as a baseline strategy. For ER_10, we use the English dataset used in GR with original responses. The Phase 1 dataset is Instruct for each row, while Phase 2 is MultiAlpaca. The first row for both Mistral-7B and LLaMA-3-8B provides numbers for Instruct-MultiAlpaca (from Table 1 &  2).
5 Mitigating Strategies for CFT